### v1

# PoolingVQ: A VQVAE Variant for Reducing Audio Redundancy and Boosting Multi-Modal Fusion in Music Emotion Analysis

Abstract
Multimodal music emotion analysis leverages audio and MIDI modalities to enhance performance. While mainstream approaches focus on complex feature extraction networks, we posit that shortening the length of audio sequence features to mitigate redundancy, especially in contrast to MIDI’s compact representation, may effectively boost task performance. To achieve this, we developed PoolingVQ by combining Vector Quantized Variational Autoencoder (VQVAE) with spatial pooling, which directly compresses audio feature sequences through local aggregation to reduce redundancy, then devised a two-stage co-attention approach to fuse audio and MIDI information. Experimental results on the public datasets EMOPIA and VGMIDI demonstrate that our multimodal framework achieves state-of-the-art overall performance, with PoolingVQ yielding some improvement.
Index Terms—  Multimodal, VQVAE, Cross-attention, Local aggregation, Sequence compression
1 Introduction
Music Emotion Analysis (MEA) models music-emotion links, enabling machines to perceive/quantify affect and supporting applications like recommendation and therapy via accurate emotional interpretation. Constrained by scarce datasets, current MEA focuses on utterance-level tasks : classification (Russell’s quadrant model, arousal/valence) and regression (predicting scores for these dimensions)—adapting to data limits while decoding emotions.
Early MEA methods used rule-based approaches with musicological features (mode, tempo, harmony [2, 3, 4]) —interpretable but limited by handcrafted features and expert dependence, restricting expressiveness and generalization. Deep learning enabled CNN/RNN-based models to capture abstract audio/lyric representations [5, 6, 7], yet most focused on single modalities, missing multi-signal emotional cues. Advances in deep learning brought multimodal architectures integrating audio, MIDI, lyrics [8, 9, 10], while pre-training yielded strong unimodal models: MusicBERT [11] (Transformer-based, symbolic music) captures structure and harmony; MERT [12] (diverse audio pre-training) provides rich acoustic representations. However, most existing methods still rely primarily on feature extraction techniques, with little consideration given to the redundancy in audio features [13, 14] and the cross-modal sequence length discrepancy between audio and MIDI features [15, 16].
In this case, we focus on sequence length adjustment through local aggregation to eliminate audio redundancy and reduce cross-modal length discrepancies. Our method is inspired by VQVAE’s local mapping capability in quantization [17] and the aggregation effects of spatial pooling methods [18], leading us to propose a novel VQVAE variant: PoolingVQ. As shown in Figure 1, this variant is applied to the audio branch and integrates pooling strategies.
Specifically, the PoolingVQ framework consists of the following two key steps: first, music audio clips are encoded by the frozen pre-trained MERT [^1] https://huggingface.co/m-a-p/MERT-v1-95M model. Next, the VQVAE codebook is initialized via K-means clustering on features from K batches; during training, a sliding window then processes the VQVAE codebook index sequence, where index variations guide optimal pooling for local features. As shown in Figure 2, the length of audio features (max 60s duration) is shortened to a third of the original; similarly, the length gap relative to MIDI features (length fixed at 512) is reduced to a third of its original magnitude. Subsequently, compressed audio is then fused with MIDI-BERT [^2] https://github.com/wazenmai/MIDI-BERT features via a two-stage cross-attention module for inter-modal interaction, then fed to a classifier for emotion prediction.
We evaluated on diverse datasets (EMOPIA: pure piano; VGMIDI: game soundtracks), showing PoolingVQ reduces audio redundancy while boosting performance. Compared to blind aggregation with only average/max pooling, its VQVAE-guided local mapping better preserves useful information, balancing redundancy reduction and feature preservation.
In summary, our contributions are threefold: 1) A novel VQVAE variant via codebook index-guided adaptive pooling for audio redundancy reduction. 2) Validation that local audio feature aggregation enhances cross-modal fusion. 3) Demonstration of PoolingVQ’s generalizability across diverse datasets (EMOPIA, VGMIDI).
Fig. 1: Architectures of our proposed model, which consists of three modules: pre-train model extractor module, music audio representation Local Aggregation module and multi-modal fusion module.
2 Method
As shown in Figure 1, our model comprises three modules: pre-trained extractor, audio Local Aggregation, and multi-modal fusion. The pre-trained extractor uses frozen MERT and MIDIBERT to extract audio and MIDI features, respectively. The PoolingVQ-based Local Aggregation processes audio: MERT features are codebook-mapped into local sequences, then these MERT features (in the form of local sequences) undergo pooling operations (max, average, or weighted average). Pooled audio and MIDIBERT MIDI features enter a fusion module with two cross-attentions. Fused features are fed to a classifier for emotion prediction.
2.1 Pre-trained Model Extractor Module
This module employs frozen pre-trained models MERT-95M and MIDIBERT to extract feature representations for audio and MIDI modalities, respectively. For the audio branch, the input audio data is processed according to the specifications of MERT-95M, with the sampling rate converted to 22400Hz. To align with the duration constraints of the modal branches and avoid random truncation of audio, the fixed length of audio data is set to 60 seconds. This setting is determined based on the duration distribution of the EMOPIA [19] and VGMIDI [20] datasets, ensuring it matches the maximum length of 512 that MIDIBERT can cover, thus maintaining consistency in temporal information across modalities.
2.2 Local Aggregation Module
The pooling compression module is centered on PoolingVQ, combining the VQVAE codebook and pooling strategies. It accepts audio representations extracted from the frozen pre-trained model MERT-95M.
In this work, VQVAE models the local feature distributions of $\mathbf{X}=\{x_{i}\in\mathbb{R}^{D},\,i=1,\ldots,T\}$, utilizing index sequences to capture local variations and determine corresponding pooling strategies. The framework involves two stages: pre-training, where K batches of data are cached for k-means clustering to generate P cluster centers as VQVAE’s initial codebook; and formal training, employing classical VQVAE mapping with precomputed similarity (higher similarity guiding optimal mapping):
$\cos(x_{i},P_{j})=\frac{x_{i}\cdot P_{j}}{\|x_{i}\|\|P_{j}\|},$ (1)
$c_{i}=\arg\max\left(\cos(x_{i},P_{1}),...,\cos(x_{i},P_{p})\right),$ (2)
where cos($x_{i}$, $P_{j}$) is the cosine similarity between local feature $x_{i}$ and the $j$-th codebook vector $P_{j}$, and $c_{i}$ denotes the index of the codebook vector assigned $x_{i}$ (i.e., the index corresponding to the maximum cosine similarity).
Pooling strategies are selected via a sliding window (size 5, stride 3) based on index sequence $\mathbf{C}=\{c_{i}\in\mathbb{N},\,i=1,\ldots,T\}$: average pooling for invariant indices, weighted average for less than or equal to 3 distinct indices, and max pooling for greater than 3 distinct indices.
$Pool(w)=\begin{cases}AvgP&|\{c_{i}\}_{w}|=1,\\ WtAvgP&1<|\{c_{i}\}_{w}|\leq 3,\\ MaxP&|\{c_{i}\}_{w}|>3,\end{cases}$ (3)
where $Pool(w)$ is the pooling strategy for window $w$; $\{c_{i}\}_{w}$ is the index set in $w$. AvgPAvgP, WtAvgPWtAvgP, MaxPMaxP denote Average, Weighted Average, Max Pooling respectively.The sequence changes are shown in Figure 2.
Fig. 2: Changes of a single sample via PoolingVQ.
To enable more accurate modeling of local features, the traditional VQVAE [17] quantization loss fine-tunes cluster center positions.
$\mathcal{L}_{1}=\|sg(X)-P_{c}\|_{2}+\|X-sg(P_{c})\|_{2},$ (4)
where $X$ is the original input, $P_{c}$ the codebook-mapped feature vector for the index sequence $C$, and $sg(\cdot)$ is the stop-gradient operation blocking gradient flow through its argument in backpropagation.
2.3 Multi-modal Fusion Module
Modal fusion matters as heterogeneous modalities (e.g., audio, text) offer complementary info, bridging single-modal gaps. Unlike simple uniform methods (concatenation, element-wise addition), cross-attention dynamically models inter-modal links via adaptive weights for relevant pairs, cutting noise by focusing on key correlations through query-key scoring and value aggregation. The cross-modal attention calculation formula is:
$Attn(Q_{\alpha},K_{\beta},V_{\beta})=softmax\left(\frac{Q_{\alpha}K_{\beta}^{T}}{\sqrt{d}}\right)V_{\beta},$ (5)
where $\alpha$, $\beta$ denote different modalities. In this paper, Building on [21, 22], we propose a two-stage cross-attention fusion module: first, compressed audio (as queries) fuses with MIDI features; then the process reverses. This bidirectional interaction yields richer fused features.
$\begin{gathered}F_{1}=Attn(Q_{a},K_{m},V_{m}),\\ F_{final}=Attn(Q_{m},K_{1},V_{1}),\end{gathered}$ (6)
where $F_{1}$ is the initial fused feature from the first-stage cross-attention (audio query with MIDI key/value), and $F_{final}$ is the final fused feature from the second-stage cross-attention (MIDI query with $F_{1}$ key/value).
The final fused feature $F_{final}$ is fed into a fully connected classification head, outputting raw prediction values $z_{j}$ for each class $j$. For the four-class task, softmax is applied to $z_{j}$ to get predicted probabilities $\hat{y}_{j}$ ($j=1,\dots,4$). Specifically, cross-entropy loss measures the discrepancy between $\hat{y}_{j}$ and ground-truth labels $y_{j}$, defined as:
$\mathcal{L}_{CE}=-\sum_{j=1}^{4}y_{j}\log(\hat{y}_{j}).$ (7)
Therefore, the final total loss of our model is:
$\mathcal{L}_{total}=\mathcal{L}_{CE}+\mathcal{L}_{1}.$ (8)
3 EXPERIMENTS
3.1 Experimental Setup
We trained and validated our model using the EMOPIA and VGMIDI datasets. Both datasets are used for a four-classification task. Among them, EMOPIA consists of pure piano music with 1087 samples, and VGMIDI contains 200 samples of game soundtracks. To ensure comparability with previous methods, we followed the official division files of the original data: EMOPIA was divided in a 7:2:1 ratio, and VGMIDI in a 6:2:2 ratio. As noted in Section 2.1, the maximum audio duration is 60s and the maximum MIDI sequence length is 512.
The K-batch buffer caches half of the training data. Training was conducted using the Adam optimizer with parameters set as $\beta_{1}$=0.9, $\beta_{2}$=0.99, $\varepsilon$=1e-6. Specifically, the classifier adopted a learning rate of 1e-4 and weight decay of 5e-5, while the quantizer used a learning rate of 5e-5 and weight decay of 1e-5. Additionally, the ReduceLROnPlateau scheduling strategy was incorporated, featuring a decay factor of 0.7, a patience of 2 epochs, and a threshold of 1e-4.
3.2 Results and Analysis
Table 1: Performance Comparison of Different Methods on EMOPIA and VGMIDI Datasets.
Method EMOPIA VGMIDI
Acc F1 Acc F1
SVM [3] 0.477 0.476 0.451 0.377
LSTM-Ann [19] 0.647 0.563 0.417 0.260
MT-MIDIGPT [9] 0.625 0.611 0.558 0.509
MT-MIDIBERT [9] 0.676 0.664 0.498 0.453
BFAM [8] 0.822 0.770 0.625 0.547
Our proposed model 0.8953 0.8954 0.5250 0.5231
For a comprehensive comparison with existing approaches, we evaluated representative methods spanning traditional, single-modal neural network, pre-trained model, and multi-modal categories, with results summarized in Table 1. Our model achieves an F1-score of 89.54% on the EMOPIA dataset and 52.31% on VGMIDI, collectively delivering state-of-the-art overall performance. Among the comparative methods, the traditional approach SVM and the MIDI-based single-modal method LSTM-Ann exhibit limited efficacy, while pre-trained models demonstrate measurable performance gains. As a multi-modal approach, BFAM stands out as the one with the most significant progress in overall performance. By integrating pre-trained musical representations with multi-modal modeling (to capture temporal characteristics and emotional dynamics), our model demonstrates strong competitiveness. This validates that combining pre-training with multi-modal fusion effectively enhances music emotion recognition performance, offering a more advantageous technical pathway for the field.
3.3 Ablation Study
1) We evaluated the key component (Local Aggregation module vq) and compared the performance of unimodal setups using ablation experiments, results in Figure 3.
Ablation on EMOPIA with different splits is stable (a): the “ALL” configuration consistently achieves the highest F1-score, and other configurations exhibit reproducible performance, which indicates consistent impacts of components across splits. When ablating on EMOPIA and the more challenging VGMIDI (b), “ALL” remains optimal, gaining significantly over single-modal setups (only midi, only audio). Moreover, “ALL” outperforms “-vq” to different extents, showing that our PoolingVQ reduces audio redundancy while enhancing performance and thus validating consistent component effectiveness across datasets.
Fig. 3: Bar chart of ablation experiment F1-scores. (a) Experiments (1) and (2) test EMOPIA’s stability under different splits. (b) Experiments (1) and (3) test stability across EMOPIA and VGMIDI.
Fig. 4: Impact of Codebook Number (P)
2) Impact of initial codebook count (parameter P): We explored P’s influence by varying it across the EMOPIA and VGMIDI datasets, analyzing effects from experimental results, with results in Figure 4.
As the X-axis endpoints, “Only AvgP” and “Only MaxP” represent extreme states with extremely small and large codebook counts, respectively. On both EMOPIA and VGMIDI datasets, performance gradually rises as codebook numbers approach intermediate values. This indicates that, compared to a single and arbitrary aggregation approach, the mapped sequences of VQVAE codebooks can guide the selection of optimal aggregation strategies. Additionally, the peak intervals of the two datasets overlap to some extent, indicating our PoolingVQ has certain universality across the two datasets with distinct characteristics.
4 CONCLUSION and future work
In this study, we propose PoolingVQ—a novel Vector Quantized Variational Autoencoder (VQVAE) variant that integrates a pooling strategy for local audio feature aggregation. Designed to mitigate audio feature redundancy while preserving fine-grained sequence information, PoolingVQ enhances the subsequent two-stage cross-modal fusion. Experimental results demonstrate our method reduces local redundancy in audio sequences while achieving effective multi-modal fusion improvement, and ablation experiments confirm its significant applicability across two datasets from distinct domains. In future work, we will explore using codebooks fitted to local emotional information for local aggregation, an approach that better aligns with the continuous nature of emotional shifts in music. Additionally, our method may provide a redundancy removal step to facilitate the long-term modeling of audio.
5 ACKNOWLEDGEMENT
This work was supported by Hubei Provincial Key Laboratory of Metallurgical Industry Process System Science (Y202105) and High Performance Computing Center of Wuhan University of Science and Technology. We thank the reviewers for their contributions.
[…]
thebibliography